# Development of Neural Network-Based Optimal Control Pulse Generator for Quantum Logic Gates Using the GRAPE Algorithm in NMR Quantum Computer

### v2

@@ II Hamiltonian of the system and form of the control pulses
where $B_{0}$ is the static magnetic field, and $B_{1}(t)$ and $\phi(t)$ represent the amplitude and phase of the applied RF field, respectively. $\gamma_{i}$ is the gyromagnetic ratio for each nucleus in the system. Since $B_{1}(t)$ is significantly smaller than $B_{0}$, and the system is homonuclear, we use the approximation $\gamma_{i}B_{1}(t)=\omega_{1}(t)$ for any ii. The second term in $H_{\text{free}}$ represents the J-coupling between different nuclei [3].
To implement the desired quantum logic gates, we must determine $B_{1}(t)$ and $\phi(t)$. As a common technique in the GRAPE algorithm, these pulses are approximated as constant within NN discrete time steps. Within each step, the amplitude and phase remain constant, allowing us to compute the system’s overall operation by multiplying the operations from each time step. Since the Hamiltonian is constant within each step, we can readily calculate the dynamics for that period. Consequently, the total operation resulting from the pulse sequence is obtained by the product of the dynamics across all steps [43],
@@
where $u_{i}$ represents the unitary operation applied in step ii. In this algorithm, the derivative of fidelity with respect to control parameters is computed. By iteratively updating these parameters and following the gradient direction, we can converge on the optimal solution [15]. Here, we focus solely on implementing a single-qubit gate. The objective is for the overall system dynamics to produce the desired operation on the target qubit, while keeping all other qubits unaffected—i.e., they should undergo no evolution, that is:
@@
where $U_{1}$ represents the effective evolution of qubit 1 after NN steps, and $I_{j}$ describes the effective evolution of qubit $j$, ensuring that its state returns to the initial state—i.e., no evolution occurs for that qubit. Our demand is that $U_{total}$ approaches the $U_{target}$ as closely as possible.
@@ IV Testing the neural network
In Table.1, you’$l$ find information on the generated pulses. As illustrated, this method is significantly faster than the standard GRAPE approach, which typically takes at least 15 seconds. Consequently, this method is approximately three orders of magnitude faster, and can be integrated as part of the compiler in quantum computer architectures.
@@ V Experimental Test
By calculating the output NMR signal and its Fourier transform, as detailed in Appendix A, we find that the quantities ImIm, ReRe and PhasePhase for the state given in Eq. 12, evaluated around the frequency corresponding to qubit 1, yield the following results:

### v3

@@ II Hamiltonian of the system and form of the control pulses
where $B_{0}$ is the static magnetic field, and $B_{1}(t)$ and $\phi(t)$ represent the amplitude and phase of the applied RF field, respectively. $\gamma_{i}$ is the gyromagnetic ratio for each nucleus in the system. Since $B_{1}(t)$ is significantly smaller than $B_{0}$, and the system is homonuclear, we use the approximation $\gamma_{i}B_{1}(t)=\omega_{1}(t)$ for any $i$. The second term in $H_{\text{free}}$ represents the J-coupling between different nuclei [3].
To implement the desired quantum logic gates, we must determine $B_{1}(t)$ and $\phi(t)$. As a common technique in the GRAPE algorithm, these pulses are approximated as constant within $N$ discrete time steps. Within each step, the amplitude and phase remain constant, allowing us to compute the system’s overall operation by multiplying the operations from each time step. Since the Hamiltonian is constant within each step, we can readily calculate the dynamics for that period. Consequently, the total operation resulting from the pulse sequence is obtained by the product of the dynamics across all steps [43],
@@
where $u_{i}$ represents the unitary operation applied in step $i$. In this algorithm, the derivative of fidelity with respect to control parameters is computed. By iteratively updating these parameters and following the gradient direction, we can converge on the optimal solution [15]. Here, we focus solely on implementing a single-qubit gate. The objective is for the overall system dynamics to produce the desired operation on the target qubit, while keeping all other qubits unaffected—i.e., they should undergo no evolution, that is:
@@
where $U_{1}$ represents the effective evolution of qubit 1 after $N$ steps, and $I_{j}$ describes the effective evolution of qubit $j$, ensuring that its state returns to the initial state—i.e., no evolution occurs for that qubit. Our demand is that $U_{total}$ approaches the $U_{target}$ as closely as possible.
@@ IV Testing the neural network
In Table.1, you’ll find information on the generated pulses. As illustrated, this method is significantly faster than the standard GRAPE approach, which typically takes at least 15 seconds. Consequently, this method is approximately three orders of magnitude faster, and can be integrated as part of the compiler in quantum computer architectures.
@@ V Experimental Test
By calculating the output NMR signal and its Fourier transform, as detailed in Appendix A, we find that the quantities $Im$, ReRe and $Phase$ for the state given in Eq. 12, evaluated around the frequency corresponding to qubit 1, yield the following results: